# Strong time-periodic solutions to a multilayered fluid-structure interaction problem

### v1

@@ 4.1. Preliminary considerations
This subsection is of preparatory character. More precisely, we recall properties of the Stokes operator, and we take into account the block operator matrix associated with the damped plate block.
@@
Let us start with the Stokes operator. To this end, we introduce the space of solenoidal vector fields $\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})$ defined by $\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})\coloneqq\overline{\{f\in\mathrm{C}_{\mathrm{c}}^{\infty}(\Omega_{\mathrm{f}})^{3}:\operatorname{div}f=0,\enspace\text{in}\enspace\Omega_{\mathrm{f}}\}}^{\|\cdot\|_{\mathrm{L}^{q}}}$. Additionally invoking the gradient fields $\mathrm{G}^{q}(\Omega_{\mathrm{f}})\coloneqq\{\nabla f:f\in\mathrm{L}_{\mathrm{loc}}^{q}(\overline{\Omega_{\mathrm{f}}}),\enspace\nabla f\in\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}\}$, we deduce from [56, Thm. 1.4] that $\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}$ allows for the decomposition $\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}=\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})\oplus\mathrm{G}^{q}(\Omega_{\mathrm{f}})$. The resulting projection $\mathbb{P}\colon\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}\to\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})$ is referred to as Helmholtz projection, and we define the Stokes operator $A_{0}$ with homogeneous Dirichlet boundary conditions by
@@
(4.5) $A_{0}u\coloneqq\mathbb{P}\Delta u,\enspace\text{for}\enspace u\in\mathrm{D}(A_{0})=\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{f}})^{3}\cap\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}}).$
@@
For the following result, we refer to [50, Thm. 3], and we observe that $\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})$ is a UMD space, see also [1, Sec. III.4.4 and III.4.5] for more background on this concept. Thus, the boundedness of the $\mathcal{H}^{\infty}$-calculus especially implies the maximal $\mathrm{L}^{p}$-regularity, see for instance [16, Sec. 4.4].
[…]
@@
Consider $q\in(1,\infty)$. For the Stokes operator $A_{0}$ as introduced in (4.5), it holds that $-A_{0}\in\mathcal{H}^{\infty}(\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}}))$ with angle $\phi_{-A_{0}}^{\infty}=0$, and the spectral bound $s(A_{0})\coloneqq\sup\{\operatorname{Re}\lambda:\lambda\in\sigma(A_{0})\}$ satisfies $s(A_{0})<0$. In particular, $-A_{0}$ admits maximal $\mathrm{L}^{p}$-regularity on $\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})$.
@@
(4.6) $A_{\mathrm{p}}\coloneqq\begin{pmatrix}0&\operatorname{Id}\\ -P_{\mathrm{m}}\Delta_{\mathrm{p}}^{2}&\Delta_{\mathrm{p}}\end{pmatrix},\enspace\text{where}\enspace\mathrm{D}(A_{\mathrm{p}})=\left(\mathrm{W}^{4,q}(\omega)\cap\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\right)\times\left(\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\right).$
Here $\Delta_{\mathrm{p}}^{2}$ and $\Delta_{\mathrm{p}}$ represent the $\mathrm{L}^{q}(\omega)$-realization of the in-plane bi-Laplacian $\Delta_{s}^{2}$ and Laplacian $\Delta_{s}$ as introduced in Section 1 and subject to the clamped boundary conditions. The result below can be found in [41, Thm. 4.2], see also [19, Thm. 5.1] for the underlying solvability of the damped plate problem on the spaces without spatial average zero.
[…]
Let $q\in(1,\infty)$, and recall $A_{\mathrm{p}}$ from (4.6). Then there is $\mu\geq 0$ such that $-A_{\mathrm{p}}+\mu$ is $\mathcal{R}$-sectorial on $\mathrm{X}_{0}^{\mathrm{p}}$ with angle $\phi_{-A_{\mathrm{p}}+\mu}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$. In particular, $-A_{\mathrm{p}}+\mu$ has maximal $\mathrm{L}^{p}$-regularity on $\mathrm{X}_{0}^{\mathrm{p}}$.
@@ 4.2. The thick structural layer
(4.7) $L_{0}f\coloneqq\mu_{\mathrm{s}}\operatorname{div}\bigl{(}\nabla f+(\nabla f)^{\top}\bigr{)}+\lambda_{\mathrm{s}}\operatorname{div}\bigl{(}(\operatorname{div}f)\operatorname{Id}_{3}\bigr{)},\enspace\text{for}\enspace f\in\mathrm{D}(L_{0})=\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}.$
The result below on functional analytic and spectral properties of the Lamé operator from (4.7) is somewhat classical. The assertion on the spectral bound is implied by the compactness of the resolvent of $L_{0}$ together with a simple testing argument of the eigenvalue problem in $\mathrm{L}^{2}$, since the $q$-independent spectrum of $-L_{0}$ consists of discrete eigenvalues. The second part of the assertion can be deduced from the strong normal ellipticity of the Lamé operator thanks to the assumptions on the Lamé coefficients, see also [54, Sec. 6.2] and in particular [54, Thm. 6.2.11], which yields the required ellipticity and validity of the Lopatinskii-Shapiro condition so that the assertion follows from [15, Thm. 2.3].
[…]
Let $q\in(1,\infty)$, and assume that $\mu_{\mathrm{s}}$ and $\lambda_{\mathrm{s}}$ satisfy (2.5). Then it holds that $s(L_{0})<0$ and $-L_{0}\in\mathcal{H}^{\infty}(\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3})$ with angle $\phi_{-L_{0}}^{\infty}<\nicefrac{{\pi}}{{2}}$.
@@
$E_{\mathrm{s}}\binom{d_{1}}{d_{2}}=\binom{d_{2}}{L_{0}d_{1}+\delta L_{0}d_{2}},\enspace\text{for}\enspace\binom{d_{1}}{d_{2}}\in\mathrm{D}(E_{\mathrm{s}})=\mathrm{D}(L_{0})\times\mathrm{D}(L_{0}).$
However, note that the resulting operator is not closed, see [10, Sec. 2.1].
@@
(4.8) $A_{\mathrm{s}}\coloneq\begin{pmatrix}0&\operatorname{Id}\\ L_{0}&\delta L_{0}\end{pmatrix},\enspace\text{with}\enspace\mathrm{D}(A_{\mathrm{s}})=\bigl{\{}(d_{1},d_{2})\in\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}\times\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}:d_{1}+\delta d_{2}\in\mathrm{D}(L_{0})\bigr{\}},$
@@
$A_{\mathrm{s}}\binom{d_{1}}{d_{2}}=\binom{d_{2}}{L_{0}(d_{1}+\delta d_{2})},\enspace\text{for}\enspace\binom{d_{1}}{d_{2}}\in\mathrm{D}(A_{\mathrm{s}}).$
By standard interpolation theory and $-L_{0}\in\mathcal{BIP}(\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3})$ with power angle $\theta_{-L_{0}}<\nicefrac{{\pi}}{{2}}$ by Subsection 4.2, and thanks to the inclusion $U\in\mathcal{H}^{\infty}(\mathrm{X})\subset\mathcal{BIP}(\mathrm{X})$ for $U$ with $\theta_{U}\leq\phi_{U}^{\infty}$, see e. g. [16, Sec. 4.4], we get
@@
Below, we discuss functional analytic as well as spectral properties of the operator matrix $A_{\mathrm{s}}$.
[…]
@@
    We have $-A_{\mathrm{s}}\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ with angle $\phi_{-A_{\mathrm{s}}}^{\infty}<\nicefrac{{\pi}}{{2}}$.
@@
    The operator $-A_{\mathrm{s}}$ is $\mathcal{R}$-sectorial on $\mathrm{X}_{0}^{\mathrm{s}}$ with $\mathcal{R}$-angle $\phi_{-A_{\mathrm{s}}}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$. In particular, $-A_{\mathrm{s}}$ has maximal $\mathrm{L}^{p}$-regularity on $\mathrm{X}_{0}^{\mathrm{s}}$, and $A_{\mathrm{s}}$ generates a bounded analytic semigroup $(\mathrm{e}^{tA_{\mathrm{s}}})_{t\geq 0}$ on $\mathrm{X}_{0}^{\mathrm{s}}$.
@@
The first assertion of (c) immediately follows from (b) upon recalling $U\in\mathcal{H}^{\infty}(\mathrm{X})\subset\mathcal{R}\mathcal{S}(\mathrm{X})$ with $\phi_{U}^{\mathcal{R}}\leq\phi_{U}^{\infty}$, see for example [16, Sec. 4.4]. The second assertion then is a consequence of the first one thanks to the UMD property of the space $\mathrm{X}_{0}^{\mathrm{s}}$ by $q\in(1,\infty)$ and [59, Thm. 4.2], which provides a characterization of maximal $\mathrm{L}^{p}$-regularity in terms of $\mathcal{R}$-boundedness of the resolvents. The last part of the assertion of (c) is implied by the former ones and (a) in a standard way, see e. g. [54, Sec. 3.5].
Hence, it remains to show (a) and (b). In fact, we will first prove that $-A_{\mathrm{s}}+\mu\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ for some $\mu\geq 0$ and then investigate the spectral bound. Conceptually, our proof is inspired by the considerations in [10, Sec. 2]. However, let us emphasize that in [10], the authors are merely interested in the sectoriality, while we strive for the boundedness of the $\mathcal{H}^{\infty}$-calculus here. At this stage, note that perturbation theory for the latter concept is significantly more involved than for sectoriality.
@@
$B_{\mathrm{s}}\coloneqq A_{\mathrm{s}}-\begin{pmatrix}0&0\\ 0&\delta^{-1}\operatorname{Id}\end{pmatrix}=\begin{pmatrix}0&\operatorname{Id}\\ L_{0}&\delta L_{0}-\delta^{-1}\operatorname{Id}\end{pmatrix},\enspace\text{with}\enspace\mathrm{D}(B_{\mathrm{s}})=\mathrm{D}(A_{\mathrm{s}}).$
@@
$S=\begin{pmatrix}\operatorname{Id}&0\\ \delta^{-1}&\operatorname{Id}\end{pmatrix}\enspace\text{and}\enspace S^{-1}=\begin{pmatrix}\operatorname{Id}&0\\ -\delta^{-1}&\operatorname{Id}\end{pmatrix},$

### v2

@@ 5.1. Preliminary considerations
This subsection is of preparatory character. First, we recall some concepts that will be important in the sequel such as $\mathcal{R}$-sectoriality or the boundedness of the $\mathcal{H}^{\infty}$-calculus. Moreover, we recall properties of the Stokes operator, and we take into account the block operator matrix associated with the damped plate block.
We start by making precise what it means for a Banach space $\mathrm{X}$ to be a so-called UMD space. In fact, UMD represents unconditional martingale differences, and while it is a probabilistic concept, it can be characterized in terms of the Hilbert transform, namely, $\mathrm{X}$ is a UMD space if and only if the Hilbert transform is bounded on $\mathrm{L}^{p}(\mathbb{R};\mathrm{X})$ for some $p\in(1,\infty)$. Many classical function spaces such as $\mathrm{L}^{p}$-spaces, Sobolev spaces $\mathrm{W}^{s,p}$, Bessel potential spaces $\mathrm{H}^{s,p}$ or Besov spaces $\mathrm{B}_{pq}^{s}$ have this property provided $p$, $q\in(1,\infty)$, and it carries over to closed subspaces of these spaces. We also refer to [1, Sec. III.4.4 and III.4.5] for more background on this concept.
Next, we invoke the concept of $\mathcal{R}$-sectoriality for an operator. For this purpose, we first require the notion of $\mathcal{R}$-boundedness. For Banach spaces $\mathrm{X}$, $\mathrm{Y}$, a family of operators $\mathcal{T}(\mathrm{X},\mathrm{Y})$ from $\mathrm{X}$ to $\mathrm{Y}$ is said to be $\mathcal{R}$-bounded if there exist $C>0$ and $p\in[1,\infty)$ such that for all $N\in\mathbb{N}$, $T_{j}\in\mathcal{T}$, $x_{j}\in\mathrm{X}$, and for every independent, symmetric, $\{-1,1\}$-valued random variable $\varepsilon_{j}$ on a probability space $(\Omega,\mathcal{A},\mu)$, it holds that
(5.5) $\left\|\sum_{j=1}^{N}\varepsilon_{j}T_{j}x_{j}\right\|_{\mathrm{L}^{p}(\Omega;\mathrm{Y})}\leq C\cdot\left\|\sum_{j=1}^{N}\varepsilon_{j}x_{j}\right\|_{\mathrm{L}^{p}(\Omega;\mathrm{X})}.$
The smallest $C>0$ such that (5.5) is valid is then called the $\mathcal{R}$-bound of $\mathcal{T}$, and it is denoted by $\mathcal{R}(\mathcal{T})$. Let us recall that a closed linear operator $A\colon\mathrm{D}(A)\subset\mathrm{X}\to\mathrm{X}$ is sectorial, denoted by $A\in\mathcal{S}(\mathrm{X})$, if
  1. (a)
    $\overline{\mathrm{D}(A)}=\overline{\mathrm{R}(A)}=\mathrm{X}$, i.e., the operator is densely defined and has dense range,
  2. (b)
    $(-\infty,0)\subset\rho(A)$, and
  3. (c)
    there is $M\in(0,\infty)$ with $\|t(t+A)^{-1}\|_{\mathcal{L}(\mathrm{X})}\leq M$ for all $t>0$.
For $\Sigma_{\theta}\coloneqq\{z\in\mathbb{C}\setminus\{0\}:|\arg z|<\theta\}$ denoting the sector of angle $\theta\in(0,\pi]$ in the complex plane, the spectral angle $\phi_{A}$ of a sectorial operator $A$ is then defined by
$\phi_{A}\coloneqq\inf\left\{\phi:\rho(-A)\supset\Sigma_{\pi-\phi},\kern 5.0pt\sup_{\lambda\in\Sigma_{\pi-\phi}}\|\lambda(\lambda+A)^{-1}\|<\infty\right\}.$
Note that if $A$ is a sectorial operator on a Banach space $\mathrm{X}$ with spectral angle $\phi_{A}<\nicefrac{{\pi}}{{2}}$, then A-A generates a bounded analytic semigroup of angle $\nicefrac{{\pi}}{{2}}-\phi_{A}$. A sectorial operator $A$ on a Banach space $\mathrm{X}$ is called $\mathcal{R}$-sectorial if
$\mathcal{R}_{A}(0)\coloneqq\mathcal{R}\left\{t(t+A)^{-1}:t>0\right\}<\infty,$
and the $\mathcal{R}$-angle of an $\mathcal{R}$-sectorial operator $A$ is defined by
$\phi_{A}^{\mathcal{R}}\coloneqq\inf\{\theta\in(0,\pi):\mathcal{R}_{A}(\pi-\theta)<\infty\},\kern 5.0pt\text{where}\kern 5.0pt\mathcal{R}_{A}(\theta)\coloneqq\mathcal{R}\left\{\lambda(\lambda+A)^{-1}:|\arg\lambda|\leq\theta\right\}.$
In what follows, we will denote the class of $\mathcal{R}$-sectorial operators on a Banach space $\mathrm{X}$ by $\mathcal{R}\mathcal{S}(\mathrm{X})$.
The importance of the concept of $\mathcal{R}$-sectoriality in the context of maximal $\mathrm{L}^{p}$-regularity becomes apparent in the following result on the characterization of maximal $\mathrm{L}^{p}$-regularity on UMD spaces in terms of $\mathcal{R}$-sectoriality. This result is due to Weis [67, Thm. 4.2], see also [22, Thm. 4.4].
Lemma 5.2.
Consider a UMD space $\mathrm{X}$ as well as $p\in(1,\infty)$, and let $A$ be a sectorial operator on $\mathrm{X}$ with spectral angle $\phi_{A}<\nicefrac{{\pi}}{{2}}$. Then $A$ has maximal $\mathrm{L}^{p}$-regularity as made precise in Section 5 if and only if $A$ is $\mathcal{R}$-sectorial with $\mathcal{R}$-angle $\phi_{A}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$.
Another important operator theoretic concept is the boundedness of imaginary powers. In fact, we say that a sectorial operator $A$ on a Banach spaces has bounded imaginary powers, or $A\in\mathcal{BIP}(\mathrm{X})$, if $A^{\mathrm{i}s}\in\mathcal{L}(\mathrm{X})$ holds for all $s\in\mathbb{R}$, and there is a constant $C>0$ such that $\|A^{\mathrm{i}s}\|_{\mathcal{L}(\mathrm{X})}\leq C$ for $|s|\leq 1$. The associated power angle $\theta_{A}$ is defined by $\theta_{A}\coloneqq\limsup_{|s|\to\infty}\frac{1}{|s|}\log|A^{\mathrm{i}s}|$. The boundedness of imaginary powers is closely related with fractional power spaces: Consider $A\in\mathcal{S}(\mathrm{X})$ and $\alpha\in(0,\infty)$. Then the space $\mathrm{X}_{A^{\alpha}}$ is defined by $\mathrm{X}_{A^{\alpha}}\coloneqq(\mathrm{D}(A^{\alpha}),\|\cdot\|_{\alpha})$, where $\|x\|_{\alpha}\coloneqq\|x\|+\|A^{\alpha}x\|$. For the following result, we refer, e.g., to [61, Thm. 3.3.7].
Lemma 5.3.
For $A\in\mathcal{BIP}(\mathrm{X})$, and with $\cdot,\cdot]_{\theta}$ denoting the complex interpolation functor, we have $\mathrm{X}_{A^{\theta}}\simeq[\mathrm{X},\mathrm{D}(A)]_{\theta}$.
The last concept to be introduced is the boundedness of the $\mathcal{H}^{\infty}$-calculus. We say that a sectorial operator $A$ with spectral angle $\phi_{A}\in[0,\pi)$ and $\theta\in(\phi_{A},\pi)$ admits a bounded $\mathcal{H}^{\infty}$-calculus, or $A\in\mathcal{H}^{\infty}(\mathrm{X})$, if there is a constant $C>0$ such that
$\|f(A)\|_{\mathcal{L}(\mathrm{X})}\leq C\cdot\|f\|_{\mathrm{H}^{\infty}(\Sigma_{\theta})}\kern 5.0pt\text{for all}\kern 5.0ptf\in\mathrm{H}_{0}^{\infty}(\Sigma_{\theta}),$
where $\|f\|_{\mathrm{H}^{\infty}(\Sigma_{\theta})}\coloneqq\sup\{|f(\lambda)|:|\arg\lambda|<\theta\}$, and $\mathrm{H}_{0}^{\infty}(\Sigma_{\theta})$ represents the bounded and holomorphic functions on $\Sigma_{\theta}$ such that there exist $C$, $\varepsilon>0$ so that
$|f(\lambda)|\leq C\cdot\left|\frac{\lambda}{(1+\lambda)^{2}}\right|^{\varepsilon},\kern 5.0pt\text{for}\kern 5.0pt\lambda\in\Sigma_{\theta},$
Moreover, $\phi_{A}^{\infty}$ defined by the infimum of such $\theta$ is referred to as the $\mathcal{H}^{\infty}$-angle of $A$.
The following inclusion sheds some light on the relations of the previous results. We also refer to [22, Sec. 4.4] for more details. In fact, we get
(5.6) $A\in\mathcal{H}^{\infty}(\mathrm{X})\subset\mathcal{BIP}(\mathrm{X})\subset\mathcal{R}\mathcal{S}(\mathrm{X})\subset\mathcal{S}(\mathrm{X}),\kern 5.0pt\text{with}\kern 5.0pt\phi_{A}^{\infty}\geq\theta_{A}\geq\phi_{A}^{\mathcal{R}}\geq\phi_{A}.$
With regard to (5.6) and Subsection 5.1, establishing the boundedness of the $\mathcal{H}^{\infty}$-calculus with an angle strictly less than $\nicefrac{{\pi}}{{2}}$ is especially sufficient to show the maximal $\mathrm{L}^{p}$-regularity on a UMD space $\mathrm{X}$.
We complete the first part of this preparatory subsection with the following result on the preservation of the $\mathcal{R}$-sectoriality and the boundedness of the $\mathcal{H}^{\infty}$-calculus under similarity transforms. This will be crucial for the decoupling argument used in Subsection 5.4. For the assertion of (a), we refer to [22, Sec. 4.1], while the assertion of (b) can be found in [22, Prop. 2.11(vi)].
Lemma 5.4.
Consider Banach spaces $\mathrm{X}$ and $\mathrm{Y}$ as well as $T\in\mathcal{L}(\mathrm{X},\mathrm{Y})$. Let $A$ be a sectorial operator on $\mathrm{X}$, and define $A_{1}\coloneqq TAT^{-1}$.
  1. (a)
    It holds that $A\in\mathcal{R}\mathcal{S}(\mathrm{X})$ if and only if $A_{1}\in\mathcal{R}\mathcal{S}(\mathrm{Y})$ and $\phi_{A}^{\mathcal{R}}=\phi_{A_{1}}^{\mathcal{R}}$.
  2. (b)
    We have $A\in\mathcal{H}^{\infty}(\mathrm{X})$ if and only if $A_{1}\in\mathcal{H}^{\infty}(\mathrm{Y})$ and $\phi_{A}^{\infty}=\phi_{A_{1}}^{\infty}$.
@@
Let us proceed with the Stokes operator. To this end, we introduce the space of solenoidal vector fields $\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})$ defined by $\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})\coloneqq\overline{\{f\in\mathrm{C}_{\mathrm{c}}^{\infty}(\Omega_{\mathrm{f}})^{3}:\operatorname{div}f=0,\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{f}}\}}^{\|\cdot\|_{\mathrm{L}^{q}}}$. Additionally invoking the gradient fields $\mathrm{G}^{q}(\Omega_{\mathrm{f}})\coloneqq\{\nabla f:f\in\mathrm{L}_{\mathrm{loc}}^{q}(\overline{\Omega_{\mathrm{f}}}),\kern 5.0pt\nabla f\in\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}\}$, we deduce from [63, Thm. 1.4] that $\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}$ allows for the decomposition $\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}=\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})\oplus\mathrm{G}^{q}(\Omega_{\mathrm{f}})$. The resulting projection $\mathbb{P}\colon\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3}\to\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})$ is referred to as Helmholtz projection, and we define the Stokes operator $A_{0}$ with homogeneous Dirichlet boundary conditions by
@@
(5.7) $A_{0}u\coloneqq\mathbb{P}\Delta u,\kern 5.0pt\text{for}\kern 5.0ptu\in\mathrm{D}(A_{0})=\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{f}})^{3}\cap\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}}).$
@@
For the following result, we refer to [57, Thm. 3], and we observe that $\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})$ is a UMD space. Thus, the boundedness of the $\mathcal{H}^{\infty}$-calculus especially implies the maximal $\mathrm{L}^{p}$-regularity by (5.6) together with Subsection 5.1.
Lemma 5.5.
@@
Consider $q\in(1,\infty)$. For the Stokes operator $A_{0}$ as introduced in (5.7), it holds that $A_{0}\in\mathcal{H}^{\infty}(\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}}))$ with angle $\phi_{-A_{0}}^{\infty}=0$, and the spectral bound $s(A_{0})\coloneqq\sup\{\operatorname{Re}\lambda:\lambda\in\sigma(A_{0})\}$ satisfies $s(A_{0})<0$. In particular, $A_{0}$ admits maximal $\mathrm{L}^{p}$-regularity on $\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}})$.
@@
(5.8) $A_{\mathrm{p}}\coloneqq\begin{pmatrix}0&\operatorname{Id}\\ -P_{\mathrm{m}}\Delta_{\mathrm{p}}^{2}&\Delta_{\mathrm{p}}\end{pmatrix},\kern 5.0pt\text{where}\kern 5.0pt\mathrm{D}(A_{\mathrm{p}})=\left(\mathrm{W}^{4,q}(\omega)\cap\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\right)\times\left(\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega)\right).$
Here $\Delta_{\mathrm{p}}^{2}$ and $\Delta_{\mathrm{p}}$ represent the $\mathrm{L}^{q}(\omega)$-realization of the in-plane bi-Laplacian $\Delta_{s}^{2}$ and Laplacian $\Delta_{s}$ as introduced in Section 1 and subject to the clamped boundary conditions. The result below can be found in [48, Thm. 4.2], see also [25, Thm. 5.1] for the underlying solvability of the damped plate problem on the spaces without spatial average zero.
Lemma 5.6.
Let $q\in(1,\infty)$, and recall $A_{\mathrm{p}}$ from (5.8). Then there is $\mu\geq 0$ such that $A_{\mathrm{p}}+\mu$ is $\mathcal{R}$-sectorial on $\mathrm{X}_{0}^{\mathrm{p}}$ with angle $\phi_{-A_{\mathrm{p}}+\mu}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$. In particular, $A_{\mathrm{p}}+\mu$ has maximal $\mathrm{L}^{p}$-regularity on $\mathrm{X}_{0}^{\mathrm{p}}$.
@@ 5.2. The thick structural layer
(5.9) $L_{0}f\coloneqq\mu_{\mathrm{s}}\operatorname{div}\bigl(\nabla f+(\nabla f)^{\top}\bigr)+\lambda_{\mathrm{s}}\operatorname{div}\bigl((\operatorname{div}f)\operatorname{Id}_{3}\bigr),\kern 5.0pt\text{for}\kern 5.0ptf\in\mathrm{D}(L_{0})=\mathrm{W}^{2,q}(\Omega_{\mathrm{s}})^{3}\cap\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}.$
The result below on functional analytic and spectral properties of the Lamé operator from (5.9) is somewhat classical. The assertion on the spectral bound is implied by the compactness of the resolvent of $L_{0}$ together with a simple testing argument of the eigenvalue problem in $\mathrm{L}^{2}$, since the $q$-independent spectrum of $L_{0}$ consists of discrete eigenvalues. The second part of the assertion can be deduced from the strong normal ellipticity of the Lamé operator thanks to the assumptions on the Lamé coefficients, see also [61, Sec. 6.2] and in particular [61, Thm. 6.2.11], which yields the required ellipticity and validity of the Lopatinskii-Shapiro condition so that the assertion follows from [21, Thm. 2.3].
Lemma 5.7.
Let $q\in(1,\infty)$, and assume that $\mu_{\mathrm{s}}$ and $\lambda_{\mathrm{s}}$ satisfy (3.5). Then it holds that $s(L_{0})<0$ and $L_{0}\in\mathcal{H}^{\infty}(\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3})$ with angle $\phi_{-L_{0}}^{\infty}<\nicefrac{{\pi}}{{2}}$.
@@
$E_{\mathrm{s}}\binom{d_{1}}{d_{2}}=\binom{d_{2}}{L_{0}d_{1}+\delta L_{0}d_{2}},\kern 5.0pt\text{for}\kern 5.0pt\binom{d_{1}}{d_{2}}\in\mathrm{D}(E_{\mathrm{s}})=\mathrm{D}(L_{0})\times\mathrm{D}(L_{0}).$
However, note that the resulting operator is not closed, see [16, Sec. 2.1].
@@
(5.10) $A_{\mathrm{s}}\coloneq\begin{pmatrix}0&\operatorname{Id}\\ L_{0}&\delta L_{0}\end{pmatrix},\kern 5.0pt\text{with}\kern 5.0pt\mathrm{D}(A_{\mathrm{s}})=\bigl\{(d_{1},d_{2})\in\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}\times\mathrm{W}_{0}^{1,q}(\Omega_{\mathrm{s}})^{3}:d_{1}+\delta d_{2}\in\mathrm{D}(L_{0})\bigr\},$
@@
$A_{\mathrm{s}}\binom{d_{1}}{d_{2}}=\binom{d_{2}}{L_{0}(d_{1}+\delta d_{2})},\kern 5.0pt\text{for}\kern 5.0pt\binom{d_{1}}{d_{2}}\in\mathrm{D}(A_{\mathrm{s}}).$
By Subsection 5.2 as well as (5.6), we especially have $L_{0}\in\mathcal{BIP}(\mathrm{L}^{q}(\Omega_{\mathrm{s}})^{3})$ with power angle $\theta_{-L_{0}}<\nicefrac{{\pi}}{{2}}$, so by Subsection 5.1, we get
@@
Below, we discuss functional analytic as well as spectral properties of the operator matrix $A_{\mathrm{s}}$. Note that the assumption $\delta>0$ is indispensable for obtaining the following result on the boundedness of the $\mathcal{H}^{\infty}$-calculus and thus also the maximal regularity of the operator matrix associated with the thick structural layer.
Proposition 5.9.
@@
    We have $A_{\mathrm{s}}\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ with angle $\phi_{-A_{\mathrm{s}}}^{\infty}<\nicefrac{{\pi}}{{2}}$.
@@
    The operator $A_{\mathrm{s}}$ is $\mathcal{R}$-sectorial on $\mathrm{X}_{0}^{\mathrm{s}}$ with $\mathcal{R}$-angle $\phi_{-A_{\mathrm{s}}}^{\mathcal{R}}<\nicefrac{{\pi}}{{2}}$. In particular, $A_{\mathrm{s}}$ has maximal $\mathrm{L}^{p}$-regularity on $\mathrm{X}_{0}^{\mathrm{s}}$, and $A_{\mathrm{s}}$ generates a bounded analytic semigroup $(\mathrm{e}^{tA_{\mathrm{s}}})_{t\geq 0}$ on $\mathrm{X}_{0}^{\mathrm{s}}$.
@@
The first assertion of (c) immediately follows from (b) upon recalling (5.6). The second assertion then is a consequence of the first one thanks to the UMD property of the space $\mathrm{X}_{0}^{\mathrm{s}}$ by $q\in(1,\infty)$ and the characterization of maximal $\mathrm{L}^{p}$-regularity in terms of $\mathcal{R}$-sectoriality. The last part of the assertion of (c) is implied by the former ones and (a) in a standard way, see, e.g., [61, Sec. 3.5].
Hence, it remains to show (a) and (b). In fact, we will first prove that $A_{\mathrm{s}}+\mu\in\mathcal{H}^{\infty}(\mathrm{X}_{0}^{\mathrm{s}})$ for some $\mu\geq 0$ and then investigate the spectral bound. Conceptually, our proof is inspired by the considerations in [16, Sec. 2]. However, let us emphasize that in [16], the authors are merely interested in the sectoriality, while we strive for the boundedness of the $\mathcal{H}^{\infty}$-calculus here. At this stage, note that perturbation theory for the latter concept is significantly more involved than for sectoriality.
@@
$B_{\mathrm{s}}\coloneqq A_{\mathrm{s}}-\begin{pmatrix}0&0\\ 0&\delta^{-1}\operatorname{Id}\end{pmatrix}=\begin{pmatrix}0&\operatorname{Id}\\ L_{0}&\delta L_{0}-\delta^{-1}\operatorname{Id}\end{pmatrix},\kern 5.0pt\text{with}\kern 5.0pt\mathrm{D}(B_{\mathrm{s}})=\mathrm{D}(A_{\mathrm{s}}).$
@@
$S=\begin{pmatrix}\operatorname{Id}&0\\ \delta^{-1}&\operatorname{Id}\end{pmatrix}\kern 5.0pt\text{and}\kern 5.0ptS^{-1}=\begin{pmatrix}\operatorname{Id}&0\\ -\delta^{-1}&\operatorname{Id}\end{pmatrix}$,